# The global attractor of the inelastic linear Boltzmann equation in a gravity field for Maxwell molecules

### v1

@@ 1.4 Notations
  • these spaces will be equipped with the uniform norm, denoted by $||\cdot||_{\infty}$.
@@
    $||\cdot||_{\infty,p}=\sup_{\xi}\frac{|\cdot|}{1+|\xi|^{p}},$
    and the space $\mathcal{C}_{p}(\mathbb{R}^{d},\mathbb{C})$, denoted also by $\mathcal{C}_{p}$ in short, as the subspace of $\mathcal{C}(\mathbb{R}^{d})$ with a finite $||\cdot||_{\infty,p}$ norm,
  • the supremum norm in $t$ in the interval $I\subset\mathbb{R}$ for the $||\cdot||_{\infty,p}$ norm: $|||\cdot|||_{I}:=\sup_{t\in I}||\cdot||_{\infty,p}$,
  • the total variation norm: $\|\mu\|_{\mathscr{M}}$,
[…]
@@
W recall the usual embeddings of the functional spaces that we will constantly use throughout this article:
$\mathcal{C}^{\infty}_{c}(\mathbb{R}^{d})\subseteq\overline{\mathcal{C}^{\infty}_{c}(\mathbb{R}^{d})}^{||\cdot||_{\infty}}=\mathcal{C}_{0}(\mathbb{R}^{d})\subseteq\mathcal{C}_{b}(\mathbb{R}^{d}).$ (1.24)
For a measure $\mu\in\mathscr{M}(X)$ we denote its total variation norm, or equivalently its norm as a continuous linear functional on $\mathcal{C}_{0}(X)$, by $\|\mu\|_{\mathscr{M}}$. Moreover we notice that since the measure is finite, the equality $\|\mu\|_{\mathscr{M}}=|\mu|(X)$ is well-defined. We will use the weak-*- ∗ topology on the space $\mathcal{C}_{0}(X)$, and we will refer to this topology as the weak topology of $\mathscr{M}(X)$, as customary in the literature (for instance, in [36]).
@@ 1.5 Main results
$\begin{cases}\displaystyle\partial_{t}f(t,v)+a\cdot\partial_{v}f(t,v)=\int_{\mathbb{S}^{d-1}}B(|\hskip 1.42262pt^{\prime}N\cdot\omega|,|\hskip 1.42262pt^{\prime}v|)\frac{1}{|Jc(^{\prime}v)|}f(t,\hskip 0.28453pt^{\prime}v)d\omega-f(t,v)\int_{\mathbb{S}^{d-1}}B(|N\cdot\omega|,|v|)d\omega,&\\ f(0,v)=f_{0}(v).\end{cases}$ (1.25)
Using the form of the Maxwell molecule collision kernel introduced in (1.1), i.e. $B(|N\cdot\omega|,|v|)=b(|N\cdot\omega|)$, as well as (1.10), (1.25) then becomes
$\begin{cases}\displaystyle\partial_{t}f(t,v)+a\cdot\partial_{v}f(t,v)=\int_{\mathbb{S}^{d-1}}\frac{b(|\hskip 1.42262pt^{\prime}N\cdot\omega|)}{r}f(t,^{\prime}v)\mathop{}\!\mathrm{d}\omega-\left(\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\mathop{}\!\mathrm{d}\omega\right)f,&\\ f(0,v)=f_{0}(v).\end{cases}$ (1.26)
@@
We assume that the the angular part $b$ of the collision kernel $B$ satisfies
@@
Definition 1.
Let $f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ be a non-negative finite Radon measure, let $T>0$ be a positive number and let the collision kernel $b(|N\cdot\omega|)$ satisfy Assumption 1. A measure $f\in\mathcal{C}\left([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)\right)$ is a weak solution of (1.26) if for every test function $\varphi\in\mathcal{C}^{1}\left([0,+\infty),\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})\right)$ one has
@@
Remark 1.
Notice that even if in the definition above we required the test functions to be in $\mathcal{C}^{\infty}_{c}(\mathbb{R}^{d})$ for each $t\geq 0$, equation (1.28) is well-defined as soon as the test function $\varphi$ is such that $\partial_{t}\varphi$, $\partial_{v}\varphi$ and $\varphi$ are all bounded functions.
@@
The main purpose of this paper is to prove the existence, uniqueness and stability of a stationary solution to (1.26) for Maxwell molecules interactions. To this end we introduce the following concept of a stationary solution to (1.26).
Definition 2.
@@
Theorem 1.2 (Existence and uniqueness of steady states for Maxwell molecules).
Let $d=2$ or $3$, and let $a\in\mathbb{R}^{d}$. Assume that the restitution coefficient $r$ satisfies $10<r<$ 1 (we assume in particular that $r\neq 1$). Let $b(|N\cdot\omega|)$ satisfy Assumption 1. There exists a unique stationary solution $f_{\infty}\in\mathscr{M}_{+}(\mathbb{R}^{d})$ in the sense of Definition 2. Moreover, $\{|v|,|v|^{2}\}f_{\infty}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$.
@@
Theorem 1.3 (Stability of steady states for Maxwell molecules).
Let $d=2$ or $3$, and let $a\in\mathbb{R}^{d}$. We assume that the restitution coefficient $r$ satisfies $10<r<$ 1 (we assume in particular that $r\neq 1$) and that $b(|N\cdot\omega|)$ satisfy Assumption 1. Let $f_{0}(v)\in\mathscr{M}_{+}(\mathbb{R}^{d})$ such that $|v|f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$. Let $f\in\mathcal{C}\left([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)\right)$ be the solution to (1.26) obtained in Theorem 1.1 and $f_{\infty}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ be the stationary solution to (1.26) obtained in Theorem 1.2. Then,
$f(t,v)\rightarrow f_{\infty}(v),\quad\text{as}\ \ t\to\infty,$
@@
The main tool that we will use to prove these results is the well developed machinery available for the study of the Boltzmann equations in the case of Maxwell molecules by means of the Fourier transform method that was introduced by Bobylev in [12]. In the end, it is possible to prove the local uniform convergence of the Fourier transforms of the solutions towards the Fourier transform of the steady state. As a consequence, general results of probability theory (e.g., [36]) allow to state the weak convergence of the measure-valued solutions of (1.11).
@@
[…]
@@
In particular, assuming $\mu$ to be non-negative, we can describe completely the long-time behaviour of the three first moments of the solution constructed in Theorem 1.1. This is the content of the following proposition whose proof will be provided in Section 5.
@@
Let $d=2$ or $3$, and let $a\in\mathbb{R}^{d}$. Assume that the restitution coefficient $r$ satisfies $10<r<$ 1 (we assume in particular that $r\neq 1$). Let $\mu\geq 0$ and let us assume that $f\in\mathcal{C}([0,+\infty[,\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ is a solution of (1.1) such that $\{|v|,|v|^{2}\}f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$. Then
@@ 2 Well-posedness results
To prove Theorem 1.1 we work with the adjoint equation of (1.26) relying on the duality provided by the Riesz-Markov-Kakutani theorem. We notice that for cut-off kernels it is possible to decompose $\mathscr{L}^{*}$ as
@@
$-\partial_{t}{\varphi}-a\cdot\partial_{v}\varphi=\mathscr{L}^{*}({\varphi}),$ (2.2)
@@
with $\varphi_{0}\in\mathcal{C}_{0}(\mathbb{R}^{d})$.
@@ 2.1 Markov Generator and Markov Semigroup
In order to prove the existence and uniqueness of a solution to the Cauchy problem (1.26), we will construct such a solution relying on the theory of semigroups. More precisely, we will consider (2.4) solved by the test function $\varphi$, which is the adjoint equation of (1.26). In this section, we will denote by $f$ the test functions. Such test functions solve an equation of the form:
$\partial_{t}f=T[f],$ (2.5)
@@
$f(t)=S(t)[f].$ (2.6)
First, we recall the definition of a Markov generator, which is a particular type of linear mapping acting on the space $\mathcal{C}_{0}(X)$ of continuous functions vanishing at infinity, defined on a metric space $X$ which is locally compact. This definition, as well as a general introduction to the theory of Markov generators and Markov semigroups on $\mathcal{C}_{0}$ functions defined on $X$ locally compact, can be found [47] (in particular, see Section 3). In the case of the present article, we will consider $X=\mathbb{R}^{d}$. It is worth to point out that the case of $X$ locally compact requires to develop a more sophisticated theory than in case of $X$ compact.
@@
    there exists $\lambda_{0}>0$ such that for all $\lambda\in\ ]0,\lambda_{0}]$, there exists a sequence $\left(f_{n}(\lambda)\right)_{n}\subseteq\mathscr{D}(T)$ that satisfies, denoting by $g_{n}=g_{n}(\lambda)$ the sequence $g_{n}=f_{n}-\lambda T[f_{n}]$:
    $\sup_{n\in\mathbb{N}}\|g_{n}\|_{\infty}<+\infty\hskip 14.22636pt\text{and}\hskip 14.22636ptf_{n},g_{n}\ \text{converge to}\ 1\ \text{pointwise as}\ n\rightarrow+\infty,$ (2.7)
@@
    for any $f\in\mathscr{D}(T)$ and $\lambda\geq 0$, denoting $g$ by $g=f-\lambda T[f]$, we have:
    $\inf_{x\in X}f(x)\geq\inf_{x\in X}g(x),$ (2.8)
@@
If there exists a positive $\bar{\lambda}>0$ as in the previous definition, then the image of $\text{id}-\lambda T$ is $\mathcal{C}_{0}(\Omega)$ for every $\lambda>0$.
We now turn to the definition of Markov semigroup.
@@
    $\lim_{t\rightarrow 0^{+}}\|S(t)[f]-f\|_{\infty}=0$ for every $f\in\mathcal{C}_{0}(\Omega)$,
@@
    there exists a sequence $(f_{n})_{n}\subseteq\mathcal{C}_{0}(X)$ with $\sup_{n\in\mathbb{N}}\|f_{n}\|<+\infty$ such that such $S(t)[f_{n}]\rightarrow 1$ pointwise as $n\rightarrow+\infty$ for every $t\geq 0$.
@@
Now that we have introduced the fundamental tools required for our analysis, we focus on the study of the linear operator associated with the Cauchy problem (1.26).
To this aim, we introduce the linear operator $\left(T,\mathscr{D}(T)\right)$ defined as:
$T:\mathscr{D}(T)\subseteq\mathcal{C}_{0}(\mathbb{R}^{d})\rightarrow\mathcal{C}_{0}(\mathbb{R}^{d}),\qquad T[f]:=a\cdot\partial_{v}f+\mathscr{L}^{*}[f]$ (2.9)
@@
$\mathscr{D}(T)=\Big{\{}f\in\mathcal{C}^{1}(\mathbb{R}^{d})\ :f,\>\partial_{v}f\in\mathcal{C}_{0}(\mathbb{R}^{d})\Big{\}}.$ (2.10)
We first prove that $T$ maps $\mathscr{D}(T)$ into $\mathcal{C}_{0}(\mathbb{R}^{d})$. We observe that the operator $\mathscr{L}^{*}_{+}$ maps $\mathcal{C}_{0}(\mathbb{R}^{d})$ into $\mathcal{C}_{0}(\mathbb{R}^{d})$ into itself. Indeed, since $|v^{\prime}|\geq r|v|$, we have that for any $f\in\mathcal{C}_{0}(\mathbb{R}^{d})$ and $\varepsilon>0$ there exists $M\geq 0$ such that, for any $v\in\mathbb{R}^{d}$ with $|v|\geq M$ and $\omega\in\mathbb{S}^{d-1}$, $|f(v^{\prime})|\leq\varepsilon\,.$ This implies, for $|v|\geq M$,
$\left|\mathscr{L}^{*}_{+}[f]\right|(v)\leq\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)|f(v^{\prime})|\mathop{}\!\mathrm{d}\omega\leq\varepsilon\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\mathop{}\!\mathrm{d}\omega,$
and so $\mathscr{L}^{*}_{+}[f]\in\mathcal{C}_{0}(\mathbb{R}^{d})$. Observe that we used the assumption $r\neq 0$ in a crucial manner. It is also clear that $a\cdot\partial_{v}f\in\mathcal{C}_{0}(\mathbb{R}^{d})$ and $\mathscr{L}^{*}_{-}[f]\in\mathcal{C}_{0}(\mathbb{R}^{d})$ if $f\in\mathscr{D}(T)$. As a consequence, $T$ indeed maps $\mathscr{D}(T)$ into $\mathcal{C}(\mathbb{R}^{d})$.
@@
For any positive integer $n\geq 1$, the function $f_{n}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ defined as:
$f_{n}(v)=\mathds{1}_{|v|<n}e^{-\frac{n}{n^{2}-|v|^{2}}}\forall v\in\mathbb{R}^{d},$ (2.11)
@@
  • $f_{n}$ belongs to $\mathcal{C}^{1}(\mathbb{R}^{d})$, and $f_{n}$ and its first derivative $\partial_{v}f_{n}$ are compactly supported,
  • $\sup_{n\in\mathbb{N}}\|f_{n}\|_{\infty}<+\infty$ and $\sup_{n\in\mathbb{N}}\|\partial_{v}f_{n}\|_{\infty}<+\infty$,
  • $f_{n}\rightarrow 1$ pointwise, $\partial_{v}f_{n}\rightarrow 0$ pointwise as $n\rightarrow+\infty$,
  • $f_{n}$ is decreasing, in the sense that for $v,w\in\mathbb{R}^{d}$ with $|v|\leq|w|$, we have:
    $0\leq f_{n}(w)\leq f_{n}(v)\leq 1$ (2.12)
@@
The fact that $f_{n}$ is a $\mathcal{C}^{1}$ function is a classical exercise, and the fact that $f_{n}$ and its derivative $\partial_{v}f_{n}$ are compactly supported is a consequence of the definition.
As for a uniform bound, it is clear that $\sup_{n\in\mathbb{N}}\|f_{n}\|_{\infty}=1$. Concerning its derivative, since we have:
$\partial_{v}f_{n}(v)=\mathds{1}_{|v|<n}\frac{-2nv}{\left(n^{2}-|v|^{2}\right)^{2}}e^{-\frac{n}{n^{2}-|v|^{2}}},$ (2.13)
@@
$|\partial_{v}f_{n}|(v)\leq\frac{2n|v|}{\left(n^{2}-|v|^{2}\right)^{2}}e^{-\frac{n}{n^{2}-|v|^{2}}}\leq 2\left(\frac{n}{n^{2}-|v|^{2}}\right)^{2}e^{-\frac{n}{n^{2}-|v|^{2}}}\leq 2\sup_{\rho\in\mathbb{R}}\rho^{2}e^{-\rho},$ (2.14)
so that we determined an upper bound on $\partial_{v}f_{n}$, uniform in $v$ and $n$.
@@
Therefore, $f_{n}\rightarrow 1$ and $\partial_{v}f_{n}\rightarrow 0$ pointwise as $n\rightarrow+\infty$.
Finally, the monotonocity of $f_{n}$ follows directly from the definition. ∎
Finally, for $f\in\mathscr{D}(T)$, $g\in\mathcal{C}_{0}(\mathbb{R}^{d})$ and $\lambda>0$ given, we will consider the following equation
$f-\lambda T[f]=g,$
@@
$f+\lambda\mathscr{L}_{-}^{*}[f]-\lambda a\cdot\partial_{v}f-\lambda\mathscr{L}_{+}^{*}[f]=g$

### v2

@@ 1.4 Notations
[…]
@@
    $|\hskip-1.42262pt|\cdot|\hskip-1.42262pt|_{\infty,p}=\sup_{\xi}\frac{|\cdot|}{1+|\xi|^{p}},$
    and the space $\mathcal{C}_{p}(\mathbb{R}^{d},\mathbb{C})$, denoted also by $\mathcal{C}_{p}$ in short, as the subspace of $\mathcal{C}(\mathbb{R}^{d})$ with a finite $|\hskip-1.42262pt|\cdot|\hskip-1.42262pt|_{\infty,p}$ norm,
  • the supremum norm in $t$ in the interval $I\subset\mathbb{R}$ for the $|\hskip-1.42262pt|\cdot|\hskip-1.42262pt|_{\infty,p}$ norm: $|\hskip-1.42262pt|\hskip-1.42262pt|\cdot|\hskip-1.42262pt|\hskip-1.42262pt|_{I}:=\sup_{t\in I}|\hskip-1.42262pt|\cdot|\hskip-1.42262pt|_{\infty,p}$,
  • the total variation norm: $\|\cdot\|_{\mathscr{M}}$,
  • the operator norm: $\|\cdot\|$ or $\|\cdot\|_{X}$ in case we want to emphasize the Banach space $X$.
@@
We recall the usual embeddings of the functional spaces that we will constantly use throughout this article:
$\mathcal{C}^{\infty}_{c}(\mathbb{R}^{d})\subseteq\overline{\mathcal{C}^{\infty}_{c}(\mathbb{R}^{d})}^{|\hskip-1.42262pt|\cdot|\hskip-1.42262pt|_{\infty}}=\mathcal{C}_{0}(\mathbb{R}^{d})\subseteq\mathcal{C}_{b}(\mathbb{R}^{d}).$ (1.8)
For a measure $f\in\mathscr{M}(X)$ we denote its total variation norm, or equivalently its norm as a continuous linear functional on $\mathcal{C}_{0}(X)$, by $\|f\|_{\mathscr{M}}$. Moreover we notice that since the measure is finite, the equality $\|f\|_{\mathscr{M}}=|f|(X)$ is well-defined. We will use the weak-* topology on the space $\mathcal{C}_{0}(X)$, and we will refer to this topology as the weak topology of $\mathscr{M}(X)$, as customary in the literature (e.g., see [30]).
@@ 1.5 Main results
$\begin{cases}\displaystyle\partial_{t}f(t,v)+a\cdot\partial_{v}f(t,v)=\int_{\mathbb{S}^{d-1}}\frac{b(|\hskip 1.42262pt^{\prime}\hskip-2.13394ptN\cdot\omega|)}{r}f(t,^{\prime}\hskip-2.84526ptv)\mathop{}\!\mathrm{d}\omega-\left(\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\mathop{}\!\mathrm{d}\omega\right)f(t,v),&\\ f(0,v)=f_{0}(v),\end{cases}$ (1.9)
where
$N=\frac{v}{|v|},\hskip 14.22636pt\hskip 1.42262pt^{\prime}\hskip-2.13394ptN=\frac{\hskip 1.42262pt{}^{\prime}\hskip-1.42262ptv}{|^{\prime}\hskip-1.42262ptv|},\hskip 14.22636pt\text{and}\hskip 14.22636pt^{\prime}\hskip-1.42262ptv=v-(1+1/r)\big(v\cdot\omega\big)\omega.$ (1.10)
@@
We assume that the angular part $b$ of the collision kernel $B$ satisfies
@@
Definition 1 (Weak solution of (1.9)).
Let $f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ be a non-negative finite Radon measure, let $T>0$ be a positive number and let the collision kernel $b(|N\cdot\omega|)$ satisfy Assumption 1. A measure $f\in\mathcal{C}\left([0,+\infty[,\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)\right)$ is a weak solution of (1.9) if for every test function $\varphi\in\mathcal{C}^{1}\left([0,+\infty[,\mathcal{C}^{1}_{0}(\mathbb{R}^{d})\right)$ one has
@@
The main purpose of this paper is to prove the existence, uniqueness and stability of a stationary solution to (1.9). To this end we introduce the following concept of stationary solution to (1.9).
Definition 2 (Stationary solution to (1.9)).
@@
Theorem 1.2 (Existence and uniqueness of steady states).
Let $d=2$ or $3$, and let $a\in\mathbb{R}^{d}$. Assume that the restitution coefficient $r$ satisfies $0<r<1$ (we assume in particular that $r\neq 1$). Let $b(|N\cdot\omega|)$ satisfy Assumption 1. Then, there exists a unique stationary solution $f_{\infty}\in\mathscr{M}_{+}(\mathbb{R}^{d})$ in the sense of Definition 2. Moreover, $\{|v|,|v|^{2}\}f_{\infty}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$.
@@
Theorem 1.3 (Stability of steady states).
Let $d=2$ or $3$, and let $a\in\mathbb{R}^{d}$. We assume that the restitution coefficient $r$ satisfies $0<r<1$ (we assume in particular that $r\neq 1$) and that $b(|N\cdot\omega|)$ satisfy Assumption 1. Let $f_{0}(v)\in\mathscr{M}_{+}(\mathbb{R}^{d})$ such that $|v|f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$. Let $f\in\mathcal{C}\left([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)\right)$ be the solution to (1.9) obtained in Theorem 1.1 and $f_{\infty}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ be the stationary solution to (1.9) obtained in Theorem 1.2. Then:
$f(t,v)\rightarrow f_{\infty}(v),\quad\text{as}\ \ t\to\infty,$ (1.15)
@@
The main tool that we will use to prove Theorem 1.3 is the well developed machinery available for the study of the Boltzmann equations in the case of Maxwell molecules by means of the Fourier transform method that was introduced by Bobylev in [11]. In the end, it is possible to prove the local uniform convergence of the Fourier transforms of the solutions towards the Fourier transform of the steady state. As a consequence, general results of probability theory (e.g., [30]) allow to state the weak convergence of the measure-valued solutions of (1.9).
Remark 2.
We notice that the approach used to obtain Theorem 1.3 allows also to establish the well-posedness of the inelastic linear Boltzmann equation (1.9), under an additional assumption on the first order moment. Nevertheless, the stability of the steady state is obtained under weaker assumptions as in Theorem 1.2. Therefore, Theorem 1.3 establishes that if another steady state of (1.9) exists, then its first order moment cannot be finite. For this reason, we separate and rely on these two approaches.
Moreover, it might be that the semigroup approach can also be fruitfully applied to investigate the case of more general collision kernels.
@@
$=T^{\mu}(t)\int_{\mathbb{S}^{d-1}}\frac{1}{r}b\left(\left|\hskip 1.42262pt{}^{\prime}\hskip-2.13394ptN\cdot\omega\right|\right)f(t,^{\prime}\hskip-2.84526ptv)\mathop{}\!\mathrm{d}\omega-T^{\mu}(t)\left(\int_{\mathbb{S}^{d-1}}b\left(\left|N\cdot\omega\right|\right)\mathop{}\!\mathrm{d}\omega\right)f(t,v)$ (1.16)
@@
In particular, assuming $\mu$ to be non-negative, we can describe completely the long-time behaviour of the three first moments of the solution of the rehomogeneized inelastic Boltzmann equation (1.5), assuming that such a solution exists, as well as its first moments. This is the content of the following proposition whose proof will be provided in Section 5.
@@
Let $d=2$ or $3$, and let $a\in\mathbb{R}^{d}$. Assume that the restitution coefficient $r$ satisfies $0<r<1$ (we assume in particular that $r\neq 1$). Let $\mu\geq 0$ and let us assume that $f\in\mathcal{C}([0,+\infty[,\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ is a solution of (1.1) in the sense of Definition 1 such that $\{|v|,|v|^{2}\}f(t)\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ for any $t\geq 0$. Then:
@@ 2 Well-posedness results
To prove Theorem 1.1 we work with the adjoint equation of (1.9), relying on the duality provided by the Riesz-Markov-Kakutani theorem. We notice that for cut-off kernels it is possible to decompose $\mathscr{L}^{*}$ as
@@
$\partial_{t}{\varphi}-a\cdot\partial_{v}\varphi=\mathscr{L}^{*}({\varphi}),$ (2.2)
@@
with $\varphi_{0}\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})$.
@@ 2.1 Markov Generator and Markov Semigroup
In order to prove the existence and uniqueness of a solution to the Cauchy problem (1.9), we will construct such a solution relying on the theory of semigroups. More precisely, we will consider (2.4) solved by the test function $\varphi$, which is the adjoint equation of (1.9), and which has the form:
$\partial_{t}\varphi=T[\varphi],$ (2.5)
@@
$\varphi(t)=S(t)[\varphi_{0}],$ (2.6)
with $\varphi_{0}=\varphi(0)$.
First, we recall the definition of a Markov generator, which is a particular type of linear mappings acting on the space $\mathcal{C}_{0}(X)$ of continuous functions vanishing at infinity, defined on a metric space $X$ which is locally compact. This definition, as well as a general introduction to the theory of Markov generators and Markov semigroups on $\mathcal{C}_{0}$ functions defined on $X$ locally compact, can be found [40] (in particular, see Section 3). In the case of the present article, we will consider $X=\mathbb{R}^{d}$. It is worth to point out that the case of $X$ locally compact requires to develop a more sophisticated theory than in case of $X$ compact.
@@
    there exists $\lambda_{0}>0$ such that for all $\lambda\in\ ]0,\lambda_{0}]$, there exists a sequence $\left(\varphi_{n}(\lambda)\right)_{n}\subseteq\mathscr{D}(T)$ that satisfies, denoting by $\psi_{n}=\psi_{n}(\lambda)$ the sequence $\psi_{n}=\varphi_{n}-\lambda T[\varphi_{n}]$:
    $\sup_{n\in\mathbb{N}}\|\psi_{n}\|_{\infty}<+\infty\hskip 14.22636pt\text{and}\hskip 14.22636pt\varphi_{n},\psi_{n}\ \text{converge to}\ 1\ \text{pointwise as}\ n\rightarrow+\infty,$ (2.7)
@@
    for any $\varphi\in\mathscr{D}(T)$ and $\lambda\geq 0$, denoting $\psi$ by $\psi=\varphi-\lambda T[\varphi]$, we have
    $\inf_{x\in X}\varphi(x)\geq\inf_{x\in X}\psi(x),$ (2.8)
@@
If there exists a positive $\bar{\lambda}>0$ as in the fourth point of the previous definition, then the image of $\text{id}-\lambda T$ is equal to $\mathcal{C}_{0}(\Omega)$ for every $\lambda>0$.
We now turn to the definition of a Markov semigroup.
@@
    $\lim_{t\rightarrow 0^{+}}\|S(t)[\varphi]-\varphi\|_{\infty}=0$ for every $\varphi\in\mathcal{C}_{0}(\Omega)$,
@@
    there exists a sequence $(\varphi_{n})_{n}\subseteq\mathcal{C}_{0}(X)$ with $\sup_{n\in\mathbb{N}}\|\varphi_{n}\|<+\infty$ such that such $S(t)[\varphi_{n}]\rightarrow 1$ pointwise as $n\rightarrow+\infty$ for every $t\geq 0$.
@@
Now that we have introduced the fundamental tools required for our analysis, we focus on the study of the linear operator associated to the Cauchy problem (1.9). To this aim, we introduce the linear operator $\left(T,\mathscr{D}(T)\right)$ defined as:
$T:\mathscr{D}(T)\subseteq\mathcal{C}_{0}(\mathbb{R}^{d})\rightarrow\mathcal{C}_{0}(\mathbb{R}^{d}),\qquad T[\varphi]:=a\cdot\partial_{v}\varphi+\mathscr{L}^{*}[\varphi]$ (2.9)
@@
$\mathscr{D}(T)=\Big\{\varphi\in\mathcal{C}^{1}(\mathbb{R}^{d})\ /\ \varphi,\,\partial_{v}\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})\Big\}.$ (2.10)
We first prove that $T$ maps $\mathscr{D}(T)$ into $\mathcal{C}_{0}(\mathbb{R}^{d})$. We observe that the operator $\mathscr{L}^{*}_{+}$ maps $\mathcal{C}_{0}(\mathbb{R}^{d})$ into $\mathcal{C}_{0}(\mathbb{R}^{d})$ into itself. Indeed, since $|v^{\prime}|\geq r|v|$, for any $\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ and $\varepsilon>0$ there exists $M\geq 0$ such that, for any $v\in\mathbb{R}^{d}$ with $|v|\geq M$ and $\omega\in\mathbb{S}^{d-1}$, $|\varphi(v^{\prime})|\leq\varepsilon\,.$ This implies, for $|v|\geq M$,
$\left|\mathscr{L}^{*}_{+}[\varphi]\right|(v)\leq\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)|\varphi(v^{\prime})|\mathop{}\!\mathrm{d}\omega\leq\varepsilon\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\mathop{}\!\mathrm{d}\omega,$
and so $\mathscr{L}^{*}_{+}[\varphi]\in\mathcal{C}_{0}(\mathbb{R}^{d})$. Observe that we used the assumption $r\neq 0$ in a crucial manner. It is also clear that $a\cdot\partial_{v}\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ and $\mathscr{L}^{*}_{-}[\varphi]\in\mathcal{C}_{0}(\mathbb{R}^{d})$ if $\varphi\in\mathscr{D}(T)$. As a consequence, $T$ indeed maps $\mathscr{D}(T)$ into $\mathcal{C}(\mathbb{R}^{d})$.
@@
For any positive integer $n\geq 1$, the function $\varphi_{n}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ defined as:
$\varphi_{n}(v)=\mathds{1}_{|v|<n}e^{-\frac{n}{n^{2}-|v|^{2}}}\hskip 14.22636pt\forall v\in\mathbb{R}^{d},$ (2.11)
@@
  • $\varphi_{n}$ belongs to $\mathcal{C}^{1}(\mathbb{R}^{d})$, and $\varphi_{n}$ and its first derivative $\partial_{v}\varphi_{n}$ are compactly supported,
  • $\sup_{n\in\mathbb{N}}\|\varphi_{n}\|_{\infty}<+\infty$ and $\sup_{n\in\mathbb{N}}\|\partial_{v}\varphi_{n}\|_{\infty}<+\infty$,
  • $\varphi_{n}\rightarrow 1$ pointwise, $\partial_{v}\varphi_{n}\rightarrow 0$ pointwise as $n\rightarrow+\infty$,
  • $\varphi_{n}$ is decreasing, in the sense that for $v,w\in\mathbb{R}^{d}$ with $|v|\leq|w|$, we have:
    $0\leq\varphi_{n}(w)\leq\varphi_{n}(v)\leq 1.$ (2.12)
@@
The fact that $\varphi_{n}$ is a $\mathcal{C}^{1}$ function is a classical exercise, and the fact that $\varphi_{n}$ and its derivative $\partial_{v}\varphi_{n}$ are compactly supported is a consequence of the definition.
As for a uniform bound, it is clear that $\sup_{n\in\mathbb{N}}\|\varphi_{n}\|_{\infty}=1$. Concerning its derivative, since we have:
$\partial_{v}\varphi_{n}(v)=\mathds{1}_{|v|<n}\frac{-2nv}{\left(n^{2}-|v|^{2}\right)^{2}}e^{-\frac{n}{n^{2}-|v|^{2}}},$ (2.13)
@@
$|\partial_{v}\varphi_{n}|(v)\leq\frac{2n|v|}{\left(n^{2}-|v|^{2}\right)^{2}}e^{-\frac{n}{n^{2}-|v|^{2}}}\leq 2\left(\frac{n}{n^{2}-|v|^{2}}\right)^{2}e^{-\frac{n}{n^{2}-|v|^{2}}}\leq 2\sup_{\rho\in\mathbb{R}}\rho^{2}e^{-\rho},$ (2.14)
so that we determined an upper bound on $\partial_{v}\varphi_{n}$, uniform in $v$ and $n$.
@@
Therefore, $\varphi_{n}\rightarrow 1$ and $\partial_{v}\varphi_{n}\rightarrow 0$ pointwise as $n\rightarrow+\infty$.
Finally, the monotonocity of $\varphi_{n}$ follows directly from the definition. ∎
Besides, for $\varphi\in\mathscr{D}(T)$, $\psi\in\mathcal{C}_{0}(\mathbb{R}^{d})$ and $\lambda>0$ given, we will consider the following equation:
$\varphi-\lambda T[\varphi]=\psi,$
@@
$\varphi+\lambda\mathscr{L}_{-}^{*}[\varphi]-\lambda a\cdot\partial_{v}\varphi-\lambda\mathscr{L}_{+}^{*}[\varphi]=\psi$